OpenStudy (anonymous):

HELP! When a parallel beam of light, of intensity or energy flux, I, falls perpendicularly on an absorbing surface of area, A, the energy, E, incident in a time, t, is equal to I At. The momentum, p, transferred to the surface is given by p= E/c= IAt/c.

OpenStudy (anonymous):

1) Show that p= I At/c is homogeneous with respect to units. 2) Light of intensity 12 Wcm^-2 falls for 30 minutes on a mirror of area 460 cm^2 which reflects 80% of the incident light. Calculate the energy reflected from the mirror in this time. 3) In sunlight, the force on 2 sail, each of area 2km^2, is of the order of 10N each at a distance fro the sun equal to the Earth's orbital radius. Discuss the viability of using such sails to propel space vehicles. already in orbit, to reach the outer planets of the solar system.

OpenStudy (anonymous):

Hi, which bit are you stuck on ?

OpenStudy (anonymous):

All three questions, unfortunately.

OpenStudy (anonymous):

look at the first part, what is it actually asking you to do ?

OpenStudy (anonymous):

Well, it's saying that one has to show that p= I At/c is the same throughout? I understand that the units must cancel out and the correct unit for p should be obtained to prove that the formula is correct.

OpenStudy (anonymous):

yes, so you could simplify the units on the right hand side and show that they are the units of momentum

OpenStudy (anonymous):

Alright so then it'll be \[p = \frac{ Wm ^{-2} \times m ^{2} \times s }{ ms ^{-1} }\] Is that right so far?

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

But when I cancel out the units, I'm left with Wms^-1 but I read that the unit for momentum is kgms^-1.

OpenStudy (anonymous):

you need to simplify the unit for watts

OpenStudy (anonymous):

simplifying what you have so far u should be left with W s^2 m^-1

OpenStudy (anonymous):

what is the simplified units for watts?

OpenStudy (anonymous):

For energy flux, what if used Jm^-2 s^-1? I would end up with Jms^-1 and if i further simplified it (J), i would then get Ns^-1 which is a simplified unit for momentum.

OpenStudy (anonymous):

watts are the unit of power, which is joules per second